### v1

# Mass Shootings, Community Mobility, and the Relocation of Economic Activity

Abstract
Using foot traffic data for over 150,000 points of interest (POIs) near the sites of 42 mass shootings (2018-2022, U.S.), we evaluate the spatial-temporal impact of the tragic events on community mobility and relocation of economic activities. Visits to nearby POIs decrease, while farther away POIs experience increased foot traffic, implying that communities shift their activities away from the shooting sites. The impact is stronger when stronger trauma responses are expected. Our results suggest that mass shootings drive significant displacements of economic activities and can consequently lead to welfare losses due to distortions in optimal choices of time and location.
Keywords: Mass Shootings, Mobility, Relocation of Economic Activity, Foot Traffic.
JEL Codes: C23, D91, R11.
1 Introduction
Mass shootings have a profound effect on survivors, witnesses, and communities that experience them. The substantial loss of innocent lives is tragic on its own and clearly demonstrates that these events come at a great cost. However, the broader literature in economics, psychology, and beyond suggests that loss of life does not constitute the sole effect of mass shootings. Rather, it is accompanied by a variety of consequences on various fronts, including: worsening mental health for both direct victims and the surrounding community (Lowe & Galea, 2017; Sharkey & Shen, 2021; Wozniak et al., 2020); exacerbation of political polarization (Donohue et al., 2019; Luca et al., 2020; Yousaf, 2021); and decreases in school enrollment and performance (Beland & Kim, 2016; Levine & McKnight, 2024; Bharadwaj et al., 2021).
Although these tragic events are the subject of a growing number of papers, evaluations of their economic consequences are rare. To the best of our knowledge, Brodeur & Yousaf (2022) is the only published study that provides a comprehensive analysis of mass shootings from an economic perspective. Using annual data, the authors find that mass shootings reduce the number of jobs and establishments in targeted counties by about 1.8% and 1.3%, respectively. They also find detrimental effects on earnings and housing prices in the years following the shootings.
In contrast to Brodeur & Yousaf (2022), which presents year-over-year and county-level effects, our analysis is centered on a more immediate and direct economic outcome: changes in visits to locations within five miles of a shooting site, right after the incident. More specifically, we answer the following questions: Do mass shootings result in the displacement of foot traffic toward locations farther from the shooting sites? Do mobility changes differ by location characteristics, such as substitutability? And are these changes related to the perceived risk of victimization or traumatic responses to the event?
For our analysis, we use the mass shootings that occurred in the United States between 2018 and 2022, as documented by Mother Jones. This list excludes shootings from more conventionally motivated crimes, such as robbery or gang activities, and thus allows our analysis to be grounded on non-targeted and unpredictable events. To measure mobility, we use Safegraph’s “weekly patterns” dataset, which includes foot traffic data and relevant characteristics for points of interest (POIs). Then, we employ event study and Difference-in-Differences (DiD) regressions to estimate the causal impact of mass shootings on community mobility over a period spanning ten weeks before and twenty weeks after the incidents.
Our empirical findings can be summarized as follows. First, visits to nearby POIs decrease significantly after a mass shooting, while farther away POIs experience increased foot traffic, suggesting a relocation of economic activity. This pattern tends to amplify overtime, leading to a significant gap in weekly visits between regions. Second, POIs that can be substituted with relative ease due to their business nature or the presence of a large number of competitors in the region exhibit more intensive changes. POIs that are difficult to find alternatives, such as daycare facilities, show no significant effect. [^1] We discuss the welfare implications of the relocation of economic activities in Section 4.1. Even without measurable changes, welfare can still be affected. If the preference for the current location decreases due to a mass shooting but relocation is infeasible because of limited alternatives, a welfare loss may occur. Finally, the mobility changes and relocation of economic activity are stronger when the shooting occurs in a previously safer area, when it receives more media coverage, and for POIs that are more heavily visited before the incident. Combined, these results suggest that people attempt to conduct their activities away from the locations marked by these tragic events, especially under the conditions where stronger trauma responses are expected.
This paper contributes to the literature in several ways. First, it sheds light on an important driver for the long-term, macroeconomic effects of mass shootings found in other studies (e.g., Brodeur & Yousaf, 2022). After an event of this type, communities relocate economic activities away from the tragic site, likely incurring friction costs or efficiency losses due to distortions in optimal location choices. This can ultimately worsen various macro-level outcomes, including wages, employment rates, and housing prices. Second, this study contributes to the broader literature on mass shootings. Our results reveal behavioral changes in response to trauma, linking the economic consequences to those documented in studies on trauma psychology (Wozniak et al., 2020; Sharkey & Shen, 2021), the economics of crime (Braakman, 2012; Dustmann & Fasani, 2014), criminology and sociology (Garofalo, 1981; Hale, 1996), and public health (Marquet et al., 2020; Guite et al., 2006). Finally, this study finds that recreational spaces, educational facilities, grocery stores, and places for temporary accommodation are among the most impacted locations. This opens up an opportunity for future studies to use the variations in activities in these categories to identify other important relationships between economic variables.
The remainder of this paper is structured as follows. Section 2 details the data sets used in this paper, with some descriptive statistics summarizing important mobility patterns in the data. Section 3 describes our empirical strategy and regression models. Section 4 presents the results, and Section 5 concludes. The online appendices include additional empirical results based on alternative specifications.
2 Data and Descriptive Analysis
2.1 Mass shootings
This study uses mass shooting data from Mother Jones, [^2] https://www.motherjones.com a nonprofit newsroom producing investigative news on politics, criminal justice, and other fields. There are a few important differences between the Mother Jones mass shootings database and other lists of shootings used in the literature. First, elsewhere, following the FBI’s definition, a mass shooting is usually defined as a shooting incident resulting in four or more deaths (not including the perpetrator), but since 2013 Mother Jones’ reporting is based on a lower threshold of three or more deaths, following the mandate for federal investigation of mass shootings authorized by President Barack Obama. Second, Mother Jones’ list excludes shootings from more conventionally motivated crimes, such as armed robbery, gang activity, and domestic violence. Instead, the list includes incidents where the perpetrator’s motivations are unclear or involve mental health, a history of violent behavior, among others.
This selection of mass shootings, therefore, allows our analysis to avoid mixing in the broader issue of gun violence and instead focus on non-targeted and unpredictable shootings, establishing reasonable grounds to treat these events as exogenous shocks. Exogenous shocks facilitate identification and, in the context of the present study, are ex ante more likely to impact community mobility patterns by virtue of their unexpected nature.
Our analysis uses a total of 42 mass shootings that occurred between 2018 and 2022. The complete list of shootings, along with information on shooting date, number of deaths (excluding the perpetrator), and injuries, can be found in online Appendix A. The human cost of the shootings is substantial, with 279 people dead and 302 experiencing injuries. The deadliest of these events took place at a Walmart in El Paso, TX and at the Robb Elementary School in Uvalde, TX - both resulting in over 20 fatalities.
2.2 Community Mobility
The foot traffic data used in this study comes from Safegraph’s weekly patterns dataset, which is based on a collection of cellphone location data from a large list of applications that send pings - that is, requests for GPS data - to the phone in which they are installed. [^3] Papers using the Safegraph data include Chen & Rohla (2018), Chen et al. (2022), Chen & Pope (2020). Safegraph aggregates pings to produce a large panel-type dataset containing the number of weekly visits to distinct points of interest (POIs). [^4] See https://www.safegraph.com/guides/visit-attribution-white-paper for more details on Safegraph’s methodology In the dataset, a point of interest is any physical place of a non-residential nature, such as a restaurant, a museum, or a church, and it can belong to any of 180+ industry groups. In addition to tracking the number of visits to individual POIs within specified date ranges, the data include information that allows us to identify the locations and types of POIs: unique ID, latitude and longitude, location name, and North American Industry Classification System (NAICS) code.
Our main analysis uses a total of 4.2 million observations for about 153 thousand distinct POIs located within five miles of the mass shooting sites, covering a period of 30 weeks for each event - 10 weeks before the shooting and 20 weeks after. The five-mile radius was chosen based on nonparametric estimation results suggesting that the impact of mass shootings on mobility vanishes beyond five miles from the shooting site. [^5] Online Appendix B.2 contains the nonparametric results.
Six shooting events took place in more than one location. These events, known as sprees, require a special handling since our analysis relies heavily on the distance between a POI and its associated shooting site. In these cases, an overall centroid may be used to represent the crime scenes, but we have opted to include all crime scenes as separate shooting sites, including all POIs within a five mile radius of each of the crime scenes, and assigning the minimum distance-to-shooting-site in the event of overlaps. This strategy improves precision and is also more adequate in our analysis given that, in some cases, the distance between crime scenes is substantial (e.g., the Atlanta Massage Parlor Shootings’ victims were 30 miles apart). In online Appendix A, events are individually identified as being sprees or having truncated time series. [^6] The Safegraph data start in January 2018 and end in December 2022, so the mobility data are truncated for some shooting events. For instance, the Marjorie Stoneman Douglas High School shooting, which occurred on February 14th, 2018, does not have a full set of 10 weeks of data before the event, and it is thus truncated on the left. Seven out of the 42 mass shootings have such truncated time series.
2.3 Supplemental Datasets
Some additional datasets are used later in the paper to perform heterogeneity analysis. The first is a simple author-built media coverage dataset, capturing the number of event-related articles available online within one-week of each shooting. Articles are identified using broad internet search terms, such as “Boulder, Colorado Shooting” or “Tulsa, Oklahoma Shooting.” The second dataset consists of county-level, time-adjusted and population-weighted statistics on homicide rates based on data from the Centers for Disease Control and Prevention. [^7] https://www.countyhealthrankings.org (County Health Rankings and Roadmaps).
2.4 Descriptive Analysis
Figure 1: Descriptive Figures of Mobility Changes After Mass Shootings
(a) Changes in Visits to McDonalds Locations After Mercy Hospital Shooting
(b) Evolution of Aggregate Visits and Active POIs
(c) Evolution of Aggregate Visits and Active POIs - By Distance Band
*Note: Subfigures (b) and (c) plot the evolution of visits and active POIs, using the week immediately before the shooting as the baseline. Sprees and events with incomplete time series are excluded.
This section offers an introductory description of the mobility patterns observed in the data. Figure 1(a) shows the changes in weekly visits to McDonald’s locations in Chicago after the Mercy Hospital shooting on November 19th, 2018.888This exercise uses a franchise to minimize heterogeneity across locations. The y-axis represents the changes in weekly visits at each location compared to the week right before the shooting. Clearly, most McDonald’s locations experienced a decrease in weekly visits in both week 1 and week 8 after the shooting (red dots), but we do see some locations farther from the shooting site experienced an increase in visits in the latter period (green dots). This may suggest that farther away locations recover more quickly after the shooting, or even that a substitution mechanism is at play. That is, customers now prefer farther away locations rather than those in close proximity to the site of the incident.
We then compute an index measure of weekly aggregate visits for all non-truncated, non-spree shooting events as a preliminary exercise. The index takes a value of 100 for the week prior to the shooting (week -1) and is calculated either for all POIs up to 5 miles away from the shooting sites or separately by distance band. Distance bands are concentric zones surrounding a shooting site, with cutoffs determined by the quantiles of the distances between POIs and the shooting site. For instance, the first quartile distance band comprises POIs located within the bottom 25% of the distances, while the second quartile distance band includes POIs located between the 25th and 50th percentiles of the distances. Distance quantiles are computed at the shooting event level, and thus the sizes of the distance bands can differ from place to place. We prefer quantiles over raw distance (miles) to define our distance bands since it accounts for differences in POI density in the various shooting areas and thus captures “perceived distance” in a local community. [^9] A potential concern is that if POIs are highly concentrated in a specific area, the distance bands based on quantiles can be physically too close, failing to represent distinct areas in a community. Figures C.1 and C.2 in online Appendix C demonstrate that this is not a concern in our data. In these figures, we plot the relationship between raw distance and distance quantile, as well as the density of a given POI type at each distance quantile. Overall, there appears to be a linear relationship between raw distance and distance quantile, implying POIs in our data are in general evenly distributed over space. POI types are also quite uniformly distributed.
Figure 1(b)(i) displays a substantial reduction in aggregate visits beginning the week of the shooting, where recovery is not complete even after 20 weeks. Figure 1(c)(i) shows a similar decline in aggregate visits across all distance bands immediately following a mass shooting, but a faster recovery is observed for the POIs in the third and fourth quartile distance bands. This aligns with the patterns observed in Figure 1(a). Visits to the POIs in the third and fourth quartile distance bands rise even above the pre-shooting levels, suggesting a potential substitution effect.
Figures 1(b)(ii) and 1(c)(ii) plot a similar index for active POIs, which displays the evolution of the number of POIs for which nonzero visits are reported in the data. The time-evolution of this index is much more subtle, indicating that the changes in mobility patterns are neither driven by POI closures, nor merely by missing some of the POIs during the period of analysis.
3 Empirical Strategy
Our empirical strategy for identifying the causal impact of mass shootings on community mobility and relocation of economic activities hinges on the assumption that there is no fundamental change in community mobility prior to the week of mass shooting. We argue that this is a reasonable assumption given that our analysis focuses on non-targeted and unpredictable shootings, as discussed in Section 2.1. Also, the pre-trends observed in the graphs in this section do not show any notable changes near the week of mass shooting, supporting the assumption.
We use the following event study design to estimate the causal impact of mass shootings:
$\quad y_{i,t}$ $=$ $\sum_{b\in\mathcal{B}}\sum_{\ell=-10}^{\ell=-2}\beta_{b,\ell}D_{i,t}^{b,\ell}+\sum_{b\in\mathcal{B}}\sum_{\ell=0}^{\ell=20}\beta_{b,\ell}D_{i,t}^{b,\ell}+\delta_{Y(ST_{i})\times b}+\gamma_{M(ST_{i})\times b}+\epsilon_{i,t},$ (1)
where $y_{i,t}$ is weekly visits to POI $i$ in week $t$, the set $\mathcal{B}$ includes distance bands, [^10] See Section 2.4 for the definition of our distance bands. In our analysis, we use either two-quantile or four-quantile distance bands. Online Appendix B.2 includes some nonparametric estimation results, demonstrating that our results are robust to the cutoffs used to define the distance bands. and $\ell$ is the week number relative to shooting omitting the week right before the shooting (baseline week). The dummy $D_{i,t}^{b,\ell}=\textbf{1}\{B_{i}=b,t-ST_{i}=\ell\}$, where $B_{i}$ is the distance band to which POI $i$ belongs, and $ST_{i}$ is the week of the mass shooting associated with POI $i$. Therefore, $D_{i,t}^{b,\ell}$ is an indicator for POI $i$ being in distance band $b$ and relative week $\ell$, and thus $\beta_{b,\ell}$ calculates the average change in visits in distance band $b$ in week $\ell$ compared to the band’s weekly visits in the week right before the shooting. The specification includes distance band-specific fixed effects for shooting year, $\delta_{Y(ST_{i})\times b}$, and month, $\gamma_{M(ST_{i})\times b}$, controlling for potential business cycles at higher-levels. The standard errors are clustered at the POI level. [^11] We considered POI fixed effects, instead of the band-specific shooting year and month fixed effects. The parameter estimates were very similar to those reported in this paper, but caused computation issues when clustering the standard errors at the POI level.
We also compute DiD-style estimates that compare the magnitude of the impact across different distance bands (e.g., Figure 2(b)) by reparameterizing the model (1) such that
$\quad y_{i,t}$ $=$ $\sum_{\ell\in\mathcal{G}}\beta_{\ell}D_{i,t}^{\ell}+\sum_{b\in\mathcal{B}/\{b^{\ast}\}}\sum_{\ell\in\mathcal{G}}\beta_{b,\ell}D_{i,t}^{b,\ell}+\delta_{Y(ST_{i})\times b}+\gamma_{M(ST_{i})\times b}+\epsilon_{i,t},$ (2)
where $\mathcal{G}=\{-10,-9,\ldots,20\}/\{-1\}$, $D_{i,t}^{\ell}=\textbf{1}\{t-ST_{i}=\ell\}$, $b^{\ast}$ is the reference band and thus $\beta_{b,\ell}$ in the model (2) calculates the impact relative to that on the reference.
4 Results
4.1 Mobility Patterns after Mass Shooting
Figure 2: Impact of Mass shootings on Community Mobility
(a) Event-Study Estimates
(b) Difference-in-Differences Estimates
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. In subfigure (b), the reference group is the POIs above median distance from the shooting site.
Figure 2(a) plots event study estimates obtained from model (1) for two distance bands split across median distance from the shooting site. POIs near shooting sites experience statistically significant declines in weekly visits for most of the period. We do see some degree of recovery after week 5, but it does not persist in the later weeks. The pattern is robust to an alternative specification using four distance bands instead of two (online Appendix B), though the effect is less precisely estimated for the 1st quartile distance band, which may be due to the presence of authorities, visitors, or the media near the shooting sites.
POIs above the median distance, on the other hand, show an increase in visits starting a few weeks after the event. Although the POIs also exhibit a decrease in weekly visits until week 5 (statistically insignificant), visits return to the pre-event level in week 6 and remain above the baseline for the remainder of the period. Two potential factors may explain the increase in weekly visits to these POIs. First, the resumption of previously postponed economic activities by individuals who normally use those POIs. Second, the relocation of activities to further away places by individuals who, prior to the event, typically used the POIs near the shooting site. The former implies only a temporal displacement of activities, but the latter suggests both temporal and spatial displacements. Although the resumption of postponed activities can account for the recovery observed between weeks 5 and 7, it does not fully explain why the recovery goes beyond the pre-event level and the increased visits persist for more than three months. Another important observation in Figure 2(a) is that, after week 5, the zero horizontal line often passes through the middle of the gap between the two plots, indicating the net change in weekly visits across all regions is often close to zero.
Together, these patterns suggest some degree of substitution or relocation of economic activities across the regions. That is, visits lost from the POIs in close proximity to the crime scene are redistributed to the POIs far from it. [^12] This relocation mechanism is further supported by the analysis in Section 4.2, where we examine the impact by POI type and POI substitutability. This substitution effect, however, does not kick in immediately after the shooting. Instead, a few weeks go by before a significant gap is observed in the post-shooting trends of weekly visits between the regions. This may be due to the difficulty of substitution in the presence of search and switching costs (Klemperer, 1987; Wilson, 2012; Weizsäcker, 1984). Therefore, the relocation of economic activities implies friction costs or efficiency losses due to a deviation from the optimal choices, which, in turn, could affect more macro-level economic outcomes such as wage, unemployment rate, and housing price, among others (a la Brodeur & Yousaf, 2022). [^13] Compared to mere temporal displacement, the relocation of activities can result in more enduring and persistent social costs since the new location is likely sub-optimal compared to the previous one.
Note that the changes in weekly visits in Figure 2(a) are “per POI”; therefore, the aggregate impact can be substantial. Since there are over 150 thousand POIs in the data, the overall impact will be the product of the estimated change per POI and the number of POIs in a given distance band. For instance, the reduction of about 6 visits per POI in week 5 for those POIs below the median distance will amount to an overall decrease of almost half a million weekly visits. The median value of nonzero weekly visits in the dataset during the week prior to the shootings (the baseline week) is 20, and thus a decrease of 6 visits constitutes a 30% decline relative to the median.
The differential mobility patterns in the two regions persist over time, resulting in a significant gap in weekly visits. Figure 2(b) plots the DiD estimates for POIs below the median distance (i.e., nearby POIs) with the reference group being the POIs above the median distance (i.e., far away POIs). The graph thus directly quantifies the gap in weekly visits over time between the two regions and its significance. The gap becomes statistically significant after week 4 and increases to about 8 visits in week 17. The same pattern is observed from an alternative specification with four distance bands, although estimates for first-distance-quartile POIs lack precision (see online Appendix B).
Our results in this section are robust to the thresholds used to define distance bands. Online Appendix B.2 provides some nonparametric results based on splines (piece-wise polynomials), where we estimate the evolution of weekly visits as a function of “continuous” distance. The patterns are consistent with those reported in this section. The spline results also support setting 5 miles as the cutoff for sample selection, as estimates approach zero.
4.2 Evidence for Relocation of Economic Activity
In this section, we analyze the impact of mass shooting by POI type and POI substitutability. If individuals relocate their activities from POIs near the shooting site to those farther away as response to the tragic event, we would expect to see similar magnitudes of increases and decreases in weekly visits for the same type POIs across the two regions. Also, if relocation is the primary driver of the mobility changes observed in the previous section, we would expect a larger impact on the POIs that are easily substitutable.
4.2.1 Impact by POI Type
Changes in weekly visits are estimated by distance band and POI category, using two-quantile distance bands and 15 POI categories. POIs are classified based on NAICS codes.
Figure 3: Change In Weekly Visits by POI type
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. ”Near”/”Far” refers to POIs below/above median distance from the shooting location.
Figure 3 reports the changes in visits and corresponding 95% confidence intervals for various categories in the week immediately after the event (Week 0) and approximately two months after the shooting (Week 8). [^14] Online Appendix D.1 includes the results for Week 2, 4, 6. The estimates indicate that decreases in visits are immediate and quite general across a variety of POI types near the shooting sites, pointing to widespread economic costs of mass shootings. Notably, for most categories, the increase in visits to farther-away POIs closely mirrors the decrease to nearby POIs, supporting the relocation hypothesis. People start favoring POIs farther away from the shooting site.
Heterogeneity of the impact across POI types is also notable. Recreation, education, temporary accommodation, food/drink stores exhibit larger changes, although their confidence intervals are also wider. The large and significant effect on education immediately after the shootings deserves a look. Disruption of educational activities not only hampers student learning, but also negatively impacts parents’ economic activities such as reduced work hours and productivity (Garcia & Cowan, 2024). Therefore, this large effect on education suggests a substantial and broad social cost.
4.2.2 Impact by POI Substitutability
The relative size of the changes in weekly visits across different POI types in Figure 3 suggests that ease of substitution might play a key role in the observed patterns. It is relatively easy to change the recreational spaces you frequent and thus we see a significant shift in visits from nearby places to far-away places for recreation. In contrast, it is difficult to find a new day care, which accounts for the zero effect on daycares two months after the shooting.
To generalize the idea, we evaluate how the substitutability of POIs is associated with the magnitude of the impact of mass shooting. We perform two DiD regressions, similar to that in Figure 2(b) (i.e., calculating the changes in weekly visits to nearby POIs using faraway POIs as the reference group), for two different groups : high-substitutability and low-substitutability POIs. Substitutability is defined in two distinct ways, as shown in Table 4(a). The Common Categories measure of substitutability accounts for the nature of business based on our own experience but can be subjective. The Competition Level measure takes into account the level of competition in a community but may be endogenous since the number of competitors or businesses is determined by local demand or differential preference among consumers. [^15] That is, even though there are many competing locations, it may be a result of many different tastes among consumers, so substitution can still be limited. Each measure has its pros and cons, so the results in this section should be used to gain insights, rather than quantify the difference.
Figures 4(b) and 4(c) report DiD estimates based on the Common Categories and the Competition Level measures, respectively. In both cases, there are distinct patterns in the evolution of foot traffic over time, with the redistribution of visits from nearby to farther-away POIs being more pronounced for highly substitutable POIs toward the end of the period. Due to the smaller sample size for each DiD regression, these differences are not statistically significant. Nevertheless, the differential patterns are evident in these plots and the patterns are robust to an alternative specification using 4-quartile distance bands in online Appendix D, supporting substitution as the key mechanism driving the mobility changes in the data.
Figure 4: Change in Weekly Visits by Substitutability
(a) Measures of Substitutability
Measure Description High Low
Common  Categories Selecting a set of POI types based on what appears easier to substitute POIs of the following types: Food/Drink Stores, Gas Stations, Recreation, Restaurants/Bars, Retail, Temporary Accommodation Others
Competition Level Selecting POIs based on number of competitors within 5 miles from the shooting site. Competitors are identified using 4-digit NAICS codes POIs that have above-median number of competitors in the region Others
(b) Using Common Categories
(c) Using Competition Level
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. The reference group is the POIs above median distance from the shooting site.
4.3 Degree of Traumatic Stress and Magnitude of Impact
The previous findings indicate that, after a mass shooting, people relocate their activities to far-away POIs. Given that the likelihood of another mass shooting occurring at the exact same location is nearly zero, the behavioral change should be understood in light of the psychological effects mass shootings have on communities. It has been well documented in the psychology and health literatures that traumatic events take a toll not only on survivors but also on their communities. Even those not directly affected by the event can develop post-traumatic stress symptoms, which is especially likely in cases of mass violence (Galea et al., 2003) and for those individuals who are physically closer to the trauma (Furr et al., 2010; Matt & Vasquez, 2008; May & Wisco, 2016).
A variety of studies find that fear of crime leads to considerable deterioration of both mental health and physical functioning (Green et al., 2002; Guite et al., 2006; Stafford et al., 2007). Wozniak et al. (2020) report that, after a 2014 university shooting, students, faculty, and staff expressed to both willingly and unwillingly reminisce on the traumatic event, and that such reminiscing was connected to the development of post-traumatic stress symptoms. Similarly, Sharkey & Shen (2021) analyze daily survey data and find that mass shootings cause people nearby to experience substantial shifts toward negative emotions. Furthermore, findings in the Economics literature confirm the existence of mental effects of crime, and add that, when people perceive an area to be unsafe, their willingness to pay decreases, and they attempt to adapt their behavior to avoid those areas (Dustmann & Fasani, 2014; Buonanno et al., 2013; Braakman, 2012).
To examine how the degree of traumatic stress in our context influences our results, we split our sample into high- and low- traumatic stress groups based on three factors: preexisting crime rate, media coverage, and pre-event visit density. We then analyze how these factors affect the magnitude of the impact of mass shooting. Table 5(a) summarizes how we split the data and predictions based on the literature. The analyses in this section are similar to those in Figures 4(b) and 4(c), and supplemental plots are provided in online Appendix E.
Figure 5: Change in Weekly Visits by Psychological Factors
(a) Factors Related to Psychological Effects
Factor High Low Prediction Why?
Preexisting Crime Shootings occurring in counties with above median homicide rates Others Communities respond less strongly to shootings in previously dangerous areas Desensitization & Stress Inoculation: Kennedy and Ceballo (2016), Ashokan et al. (2016)
Media   Coverage Shootings receiving above-median media coverage Others Communities respond more strongly to shootings with high media coverage Perceived Risk of Victimization: Mastrorocco and Minale (2018), Velasquez et al (2020)
Baseline Visit Density POIs with above-median visits prior to the shooting Others Foot traffic changes more strongly for highly-visited places Avoidance of Public Spaces: Garofalo (1981), Marquet et al (2020), Navarrete-Hernandez et al (2023)
(b) Preexisting Crime
(c) Media Coverage
(d) Baseline Visit Density
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. The reference group is the POIs above median distance from the shooting site.
4.3.1 Preexisting Crime Rates
Two theoretical frameworks from psychology inform our predictions regarding how mass shootings may have different effects depending on the preexisting levels of crime in the affected area. The first is stress inoculation, which states that exposure to smaller instances of stress can promote resilience to future stressful episodes (Ashokan et al., 2016; Gerber et al., 2018). For example, Amir & Sol (1999) find that students in an Israeli university experienced decreased levels of distress when facing a traumatic event if they had previously experienced other traumatic events. The second, desensitization, consists of emotional numbing resulting from repeated exposure to chronic community violence (Kennedy & Ceballo, 2016; Gaylord-Harden et al., 2011; Mrug et al., 2008; Ng-Mak et al., 2004). It follows that the impact of mass shootings is expected to be greater when it occurs in a previously safer area, where the event would strike people who are less habituated to violent crime.
We match each mass shooting with the homicide rate of the county where the shooting occurred in the year prior to the event. Homicide statistics are sourced from County Health Rankings and Roadmaps [^16] https://www.countyhealthrankings.org. Then, we classify the shooting areas into two groups: high-crime areas, defined as those with above-median crime rates, and low-crime areas, including all other locations.
Figure 5(b) plots the evolution of weekly visits for high- and low-crime areas. Over time, the impact of mass shootings on mobility becomes stronger when they occur in previously safer areas. That is, when communities are less habituated to violence, their reaction is stronger, shifting activities more substantially from nearby POIs to those farther away.
4.3.2 Extent of Media Coverage
As discussed previously, fear of crime leads to the deterioration of mental health outcomes. Moreover, evidence suggests that its psychological impact is driven primarily by perceived risk, not by actual recorded crimes (Pearson & Breetzke, 2014; Braakman, 2012; Hamermesh, 1999; Hale, 1996). The media, then, would play an important role through its influence on public perception. According to News Media and Crime Perceptions: Evidence from a Natural Experiment (2018), who use a staggered introduction of digital TV in Italy as a natural experiment, reduced exposure to crime news channels causes people to decrease their concerns about crime. Similarly, Velasquez et al. (2020) find that spikes in crime reporting increase people’s perceptions about the prevalence of crime, even when actual crime rates decline. It follows that, all else equal, more extensive media coverage on mass shootings is likely associated with greater psychological distress at the community level, and thus stronger impacts on community mobility patterns.
We calculate media coverage for each shooting event by the number of online news publications within one week of the event. [^17] Number of news publications is frequently used in the literature to measure media coverage (e.g., Chordia et al., 2022). We identify these news publications through an internet search using broad terms associated with the shooting event and the nearby community. For example, for the Mercy Hospital shooting on November 19th, 2018, its media coverage is the number of publications between the day of the shooting and November 26th, 2018, identified through search terms “Chicago shooting” and “Illinois shooting.” Our measure is a reasonable proxy for media coverage given that online news or searches are the primary sources of news consumption for the average person today.
Shooting events are then divided in two groups: those receiving above-median media coverage (high coverage) and those below this threshold (low coverage). Figure 5(c) plots the evolution of weekly visits for high- and low- media coverage shootings. Consistent with our prediction, more extensively publicized events tend to result in greater shifts of visits from nearby POIs to farther-away POIs.
4.3.3 Visitor Density
Fear of crime is associated with an individual’s perception of vulnerability – that is, the self-assessed likelihood of victimization. Though such perceptions might not accurately reflect past personal experience with victimization, they can translate into economically-relevant behavioral changes, such as the avoidance of certain physical areas or situations where the perceived risk of victimization is high (Hale, 1996; Garofalo, 1981; Whitley & Prince, 2005). Furthermore, the decision to avoid a physical environment depends on the individual’s assessment of potential risks, even in absence of direct connection between the location and specific instances of crime. (Marquet et al., 2020; Navarrete-Hernandez et al., 2023).
Public spaces, or places that are open to visitors and can, thus, attract large gatherings, are likely to be assessed as increasing the risk of victimization in the minds of the communities experiencing mass shootings. Figure 5(d) plots the evolution of weekly visits for POIs that, prior to the shooting, received high or low levels of visits. Here, a POI is considered to have high baseline visits if, during the week right before the shooting, it received above-median number of visits relative to all locations in the area of the shooting. All other POIs are treated as having low baseline visits.
The results are consistent with the prediction that communities respond more intensely for the locations where larger gatherings of people are expected, likely in an effort, conscious or not, to avoid situations that are perceived as conducive to victimization. POIs with low baseline visit density exhibit little to no change in visits after a mass shooting, with very narrow confidence intervals. The median number of nonzero weekly visits prior to the shootings among all POIs in the sample is 20, suggesting that the POIs in this category are generally very small and, as a result, do not show meaningful variations in visits. [^18] This may indicate that the effects of mass shootings presented elsewhere in the paper can be seen as a lower bound, as they include POIs that are expected to show little change in visits due to the nature of the business.
5 Conclusion and Discussion
In the current reality, where the tragic recurrence of mass shootings brings the issue into the spotlight of political debate every few months, the importance of studying their impact is undeniable. The broader costs of these events go well beyond the already devastating loss of life and the psychological trauma to survivors.
Our paper discusses an important, yet unexplored, cost of mass shootings: welfare loss due to displacement of economic activities. Visits to nearby POIs decrease after a mass shooting by about 30% with respect to the baseline, while visits to farther-away POIs significantly increase. The impact is not only substantial in size but also long-lasting and pervasive across a variety of location types. Evidence suggests that there is relocation of economic activity, where individuals substitute away from POIs near the shooting site, particularly for places that are easily substitutable.
A key driver of relocation is the psychological effects of these events. The impact of mass shootings is stronger when they occur in safer areas, when they receive more attention from the media, and for POIs that previously received many visitors. Relocation of activities carries great economic implications, as it would involve friction costs and efficiency losses due to distortions in optimal choices of time and location. This could, in turn, negatively impact other economic outcomes, such as earnings and productivity.
For future research, it would be desirable to estimate the monetary costs of mass shootings to local businesses using POI-level transaction data, if available. Also, the impact identified in this paper is a causal effect driven by unpredictable and exogenous events. Recreational spaces, educational facilities, grocery stores, and temporary accommodation places appear to be among the categories most affected by mass shootings. Therefore, future studies may leverage the temporal and spatial variation of activities in these categories to uncover important relationships between other economic variables.
B.2 Nonparametric Estimation using Splines
We estimate the change in weekly visits for each week number as a function of “continuous” distance from the shooting site using a spline function (piecewise polynomial). We use a cubic spline with four knots, where the quantile rank for distance, i.e., the proportion of distances that are smaller than a given distance, serves as the running variable and the four knots are set at 0.25, 0.50, 0.75, and 1. Changes are estimated with respect to the week immediately before the shootings (the baseline week). The x-axis and y-axis of the graphs in this section are “Distance Percentile” and “Change in Weekly Visits,” respectively, and therefore, the graphs plot the change in foot traffic to each distance percentile, compared to the foot traffic in week -1. The gray areas in the plots represent 95% confidence intervals and the standard errors are clustered at the POI level.
The purpose of this exercise is two-fold. First, it shows that the impact of mass shootings on mobility vanishes beyond five miles from the shooting, supporting our choice of a five-mile radius as the cutoff for sample selection. Second, the patterns here align with those reported in Section 4, demonstrating that our results are robust to the cut-offs used to define our distance bands.
Figure B.2: Change In Weekly Visits Over Time and Distance
(a) Including Sprees
(b) Excluding Sprees
Online Appendices C Distribution of POIs
C.1 Relationship Between Raw Distance and Distance Quantiles
Figure C.1: Relationship Between Raw Distance and Distance Quantiles
C.2 Distribution of POIs by POI Category
Figure C.2: Distribution of POIs over Distance, by POI Category
Online Appendices D Impact by POI Type and POI Substitutability
D.1 Impact by POI Type
Figure D.1: Evolution of Weekly Visits By Category
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. “Near”/“Far” refers to POIs below/above median distance from the shooting location.
D.2 Impact by POI Substitutability: 4-quantile distance bands
Figure D.2-1: Change In Weekly Visits by Ease of Substitution (Using Common Categories)
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. The reference group in subfigure (i) is the POIs above the median distance. The reference group in all other subfigures is the POIs in the 4th quartile distance band.
Figure D.2-2: Change In Weekly Visits by Ease of Substitution (Using Competition Level)
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. The reference group in subfigure (i) is the POIs above the median distance. The reference group in all other subfigures is the POIs in the 4th quartile distance band.
Online Appendices E Degree of Traumatic Stress and Magnitude of Impact: 4-Quantile Distance Bands
Figure E-1: Evolution of Weekly Visits by Preexisting Crime Rate
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. The reference group in subfigure (i) is the POIs above the median distance. The reference group in all other subfigures is the POIs in the 4th quartile distance band.
Figure E-2: Evolution of Weekly Visits by Media Coverage
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. The reference group in subfigure (i) is the POIs above the median distance. The reference group in all other subfigures is the POIs in the 4th quartile distance band.
Figure E-3: Evolution of Weekly Visits by Visitor Density
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. The reference group in subfigure (i) is the POIs above the median distance. The reference group in all other subfigures is the POIs in the 4th quartile distance band.